OpenStudy (anonymous):

Solve the system by elimination. {x+5y-4z=-10 {2x-y+5z=-9 {2x-10y-5z=0

OpenStudy (texaschic101):

x + 5y - 4z = -10 -->(-2) 2x - y + 5z = -9 ----------------- -2x - 10y + 8z = 20 (result of multiplying by -2) 2x - y + 5z = -9 -----------------add -11y + 13z = 11 2x - y + 5z = -9 -->(-1) 2x - 10y - 5z = 0 --------------- -2x + y - 5z = 9 (result of multiplying by -1) 2x - 10y - 5z = 0 ---------------add -9y -10z = 9 -11y + 13z = 11 -->(9) -9y - 10z = 9 --> (-11) --------------- -99y + 117z = 99 (result of multiplying by 9) 99y + 110z = -99 (result of multiplying by -11) ----------------add 227z = 0 z = 0 now sub 0 in for z -11y + 13z = 11 -11y + 13(0) = 11 -11y = 11 y = -1 now sub 0 in for z and -1 in for y x + 5y - 4z = -10 x + 5(-1) - 4(0) = -10 x - 5 = -10 x = -10 + 5 x = -5 lets check it 2x - y + 5z = -9 2(-5) -(-1) + 5(0) = -9 -10 + 1 = -9 -9 = -9 (correct) x = -5, y = -1, and z = 0